one side of the rectangle is 3 inches shorter than the other side and the perimeter is 54 inches.What are the dimensions of the rectangle.
Answer by josmiceli(19441)

The sides are x and x - 3
P = 2x + 2(x - 3)
P = 54
54 = 2x + 2(x - 3)
54 = 2x + 2x - 6
4x = 54 + 6
4x = 60
x = 15
The sides are 15 and 12
check:
P = 2*15 + 2(15 - 3)
P = 30 + 2 * 12
P = 30 + 24
P = 54
OK